# HyperSurf: Quadruped Robot Leg Capable of Surface Recognition with GRU and Real-to-Sim Transferring

### v1

@@ III-A Hardware Architecture
For conducting experiments to determine surface types, a special experimental setup was created to ensure repeatability and uniformity of the data obtained when interacting with various surface types. The creation of the stand significantly simplified and reduced the cost of the experiment, as it eliminated the need to use a full-fledged quadruped robot. The main components of the stand (1) include: a full-size leg of a quadruped robot with three degrees of freedom, mounted on a vertical stand, control electronics, an IMU sensor, and the surface for investigation.
@@
The leg mounting system allows it to move freely along the vertical guides, ensuring stability and repeatability of movements. Following control commands from the electronic unit, the leg joints are simultaneously bended and extended, allowing the leg to perform jumps in place to a height of about 20 mm, pushing off from the surface under investigation. At the end of the leg, an IMU sensor (HWT905-RS485) is attached to collect feedback.
The leg was designed and manufactured as part of a project to create an autonomous quadruped robotic system. It consists of 3 movable joints and 3 BLDC electric motors that drive its movement. The joint in contact with the surface is equipped with a resilient silicone tip to cushion the contact.
@@ III-B HyperSurf Control System
In the control system of the HyperSurf robotic leg, a combination of ROS2 and micro-ROS technologies governing its movements. Within the ROS2 package, specific leg trajectories are defined within Python scrips encapsulated in ROS2 packages. These trajectories, created to facilitate leg movement, are translated into motor angles, shaping the physical actions of the robotic limb. The computational required for these calculations is executed on an Intel NUC computer.
@@
Micro-ROS agent is facilitating the seamless interaction between the high-level computation environment and the low-level hardware control. This agent establishes a vital link between the ROS2 ecosystem running on the computer and the microcontroller system, the /joints/Commands topic published by the ROS2 environment is subscribed to by the /STM_node, which is generating PWM signals for BLDC motors drivers.
@@
III-C Choice of Surfaces for Investigation
To reliably verify the idea of surface recognition based on feedback from the IMU sensor, a diverse set of surfaces with varying properties were assembled. The interaction of the robot leg was analyzed on four types of surfaces: a sheet of EVA polymer (A), a textured ceramic surface (B), a smooth enamel-coated ceramic tile surface (C), and carpet (D).
@@ III-D Dataset Structure
Throughout the study, a dataset comprising over 1,000,000 samples (equivalent to approximately 1,5 hours of data) was amassed. Data collection encompassed four classes, including a sheet of EVA polymer, a textured ceramic surface, a smooth enamel-coated ceramic tile surface, and carpet. During the collection process, the stand operated in automatic stepping mode, varying its frequency to ensure an adequate diversity of data.
@@ III-E GRU-based Surface Recognition
Our previous method DogSurf module utilizes a GRU neural network as the foundation. The IMU data from the accelerometer and gyroscope, is collected and appended to a queue containing 100 samples. During each iteration, a Forward Pass is conducted through the firstly Standart Scaler, then through Principal component analysis (PCA) transform, and then trough bidirectional GRU - neural network.
Figure 3: Neural network architecture
Code can be accessed by link: https://github.com/Eterwait/DogSurf. The results of DogSurf module performance of the HyperSurf compared to the State-of-the-Art are listed in I.
TABLE I: Comparison of HyperSurf with the State-of-the-art Approaches
@@
Figure 4: Confusion matrix for HyperSurf
TABLE II: Results for different surfaces for HyperSurf
@@ IV-A Preparation of the stand description file
Furthermore, to simulate various surfaces within the Gazebo environment, including tiles, rubber, and brick surfaces, meticulous specifications were defined. These surface characteristics were meticulously outlined in the URDF file, with parameters carefully tuned to replicate physical attributes. Notably, parameters such as the friction coefficient, akin to a silicone surface, damping coefficient, and stiffness coefficient, were precisely calibrated to ensure accurate simulation dynamics.
@@ IV-B Kinematics Control Implementation
Leg kinematics control was realized using symbolic and numerical computation techniques. Forward kinematics determined leg component positions, while inverse kinematics adjusted joint angles for desired end-effector positions. Symbolic computation with SymPy calculated transformation matrices and coordinates, while CasADi handled numerical optimization for joint angle adjustments. This approach enabled precise leg movement control, facilitating effective interaction with the environment and data collection from various surfaces within the Gazebo simulator.
@@ IV-C Dataset Collection
V Novelty of HyperSurf System
V-A Boost of datasets collection
In comparison to collecting datasets on a quadruped robot, utilizing a one-leg test setup could increase the speed of GRU analysis by 7.5 times on average. It was achieved because of the rapid system deployment and autonomous operation in laboratory conditions.
V-B Shortening the Team for Collecting Datasets
Reducing the required collection time and utilizing a static location allowed for fewer people to be involved in the experiments. Whereas in previous work, more than five people were required due to the lengthy data collection process, in this study, only two researchers were sufficient to collect the necessary amount of data.
V-C Transferring Data
Utilizing data obtained from an IMU sensor on the stand, empirical modeling of a corresponding surface was conducted within the Gazebo simulation environment, as depicted in Figure 5. The surface modeling results within the Gazebo simulator are illustrated in Figure 6. When comparing the angular velocity of the leg between the data collected from the stand and the simulation, a simulation data accuracy of 98% was achieved.
The GRU model trained on the stand demonstrated seamless applicability to its digital twin within the simulation environment, and conversely, without necessitating additional preparation, achieving remarkable precision.
This provides ample opportunities for utilizing our system on various robotic platforms and enables the preparation of models for surfaces for which collecting datasets is challenging.
@@ VI Conclusion and Future Work
A HyperSurf leg test setup has been developed, significantly enhancing surface recognition by assisting the DogSurf classification model, achieving a precision of slippery surface recognition up to 99.9%. It enables a reduction in the time required for data collection and minimizes the need for human involvement. Additionally, it facilitates data preparation under laboratory conditions. Furthermore, we have demonstrated that the prepared data can be easily transferred to simulations, which could greatly contribute to the utilization of this method in quadruped robotics.
In the future we aim to demonstrate how our HyperSurf surface recognition system can enhance the functionality of quadruped platforms. By utilizing feedback from the GRU and process it with DogSurf architecture, we can adapt gaits to different surfaces, adjust parameters of BLDC motors, and incorporate additional mechanics to improve both the energy consumption of the robot and its capability to traverse various terrains. Furthermore, we plan to continue developing our quadruped platform (Fig. 7) and integrate our system into it, along with all features based on surface recognition feedback. Additionally, we intend to extend our system to the area of object grasping, as leveraging information about surface properties of an object could significantly enhance the effectiveness of grasping.
Figure 7: 3D model of the HyperSurf quadruped robot.

### v2

# Artem Bazhenov, Vyacheslav Kovalev, Aleksey Fedoseev and Dzmitry Tsetserukou

@@ I Introduction
In this work, we have expanded our research by systematizing the creation of digital surface duplicates within the Gazebo simulation environment, exemplified through the replication of a floor tile surface. The current state of simulation environments commonly used by the robotics community, such as Gazebo and NVIDIA’s Isaac, lacks extensive libraries of pre-defined surface types that accurately reflect the diversity of real-world terrains. This deficiency presents a significant challenge during the Sim-to-Real deployment phase, where discrepancies between simulated and actual surfaces often lead to substantial errors in robot behavior [6]. Walking robots, particularly quadrupeds, face difficulties when transitioning from controlled simulation environments to the complexities of real-world surfaces, which can differ greatly from the standard options available in these simulations. Through the development and implementation of these digital surface duplicates, we aim to reduce the trial-and-error phase associated with surface recognition and locomotion in physical environments, thereby streamlining the path from simulation to real-world deployment.
@@ III-A Hardware Architecture
To conduct experiments to determine surface types, a special experimental setup was created to ensure repeatability and uniformity of the data obtained when interacting with various surface types. The creation of the stand significantly simplified and reduced the cost of the experiment, as it eliminated the need to use a full-fledged quadruped robot. The main components of the stand (1) include a full-size leg of a quadruped robot with three degrees of freedom, mounted on a vertical stand, control electronics, an IMU sensor, and the surface for investigation.
@@
The leg mounting system is designed to allow the leg to move smoothly along the vertical guides, ensuring both stability and consistent, repeatable movements. The vertical guide mechanism is essential for controlling the range of motion, preventing lateral deviations, and ensuring that each motion, particularly the jumps, follows the intended trajectory. The leg’s movements are regulated by the electronic control unit, which sends commands to coordinate the bending and extension of the joints. This coordinated action enables the leg to perform jumping motions, lifting off the surface by approximately 20 mm. These jumps are vital for allowing the robot to interact with the surface under investigation, gathering information without excessive force or disruption. At the end of the leg, an Inertial Measurement Unit (IMU) sensor, specifically the HWT905-RS485 model, is attached. This sensor plays a crucial role in the feedback loop by providing real-time data on the leg’s position, orientation, and movement. The IMU allows the system to monitor the performance of each jump.
The leg was designed and manufactured as part of a project focused on developing an autonomous quadruped robotic system. It consists of 3 movable joints and 3 BLDC electric motors that drive its movement. The joint in contact with the surface is equipped with a resilient silicone tip to cushion the contact.
@@ III-B HyperSurf Control System
In the control system of the HyperSurf robotic leg, a combination of ROS2 and micro-ROS technologies governing its movements. Within the ROS2 package, specific leg trajectories are defined within Python scrips encapsulated in ROS2 packages. These trajectories, created to facilitate leg movement, are translated into motor angles, shaping the physical actions of the robotic limb. The computations required for these calculations were executed on an Intel NUC computer.
@@
The micro-ROS agent facilitates the seamless interaction between the high-level computation environment and the low-level hardware control. This agent establishes a vital link between the ROS2 ecosystem running on the computer and the microcontroller system managing the hardware. The /joints/Commands topic published by the ROS2 environment is subscribed to by the /STM_node. The micro-ROS agent processes these commands and translates them into Pulse Width Modulation (PWM) signals, which are then sent to the BLDC motor drivers. By facilitating this communication, the micro-ROS agent ensures that commands from the ROS2 system are accurately and efficiently executed by the hardware, allowing for precise control of the robotic system’s movements and operations.
@@
III-C Conducting the Experiment
To verify the effectiveness of surface recognition using feedback from the IMU sensor, a diverse set of surfaces with distinct properties was assembled for testing. The robotic leg’s interaction was assessed on four different types of surfaces: an EVA polymer sheet (A), a textured ceramic surface (B), a smooth enamel-coated ceramic tile (C), and a carpet (D).
The setup ensured that the leg stand remained stationary to avoid any vibrations that might interfere with the performance of the neural network. One of the surfaces was then placed underneath the robotic leg, ensuring it stayed immobile. These jumping tests lasted between 5 to 7 minutes on each surface to gather sufficient data, repeating this process four times for each surface. The surfaces were changed and the jump frequency was varied, ranging from one jump every 0.8 - 2 seconds. Data from the IMU sensor was recorded using its API throughout the experiment. This data was subsequently processed and transformed to prepare it for the training of the neural network, ensuring that the model could accurately recognize and differentiate between the various surfaces based on the sensor feedback.
@@ III-D Dataset Structure
Throughout the study, a dataset comprising over 1,000,000 samples (equivalent to approximately 1.5 hours of data) was amassed. Data collection encompassed four classes: a sheet of EVA polymer, a textured ceramic surface, a smooth enamel-coated ceramic tile surface, and carpet. During the collection process, the stand operated in automatic stepping mode, varying its frequency to ensure an adequate diversity of data. This variation in frequency was crucial for ensuring that the dataset captured a wide range of interaction scenarios, providing the necessary diversity to improve the robustness and generalization of the subsequent analysis. The large volume of data collected from these different surfaces was critical for training and refining the surface recognition algorithms used in the study.
@@ III-E GRU-based Surface Recognition
Our previous method, DogSurf module utilizes a GRU neural network as the foundation. The IMU data from the accelerometer and gyroscope is collected and appended to a queue containing 100 samples. During each iteration, a Forward Pass is conducted through the first Standard Scaler, then through Principal Component Analysis (PCA) transform, and then through bidirectional GRU neural network.
Figure 3: Neural network architecture.
Code and other supplementary materials can be accessed by link: https://github.com/Alfasaz/HyperSurf. [15] The results of DogSurf module performance of the HyperSurf compared to the State-of-the-Art are listed in Table I.
TABLE I: Comparison of HyperSurf with the State-of-the-art Approaches.
@@
Figure 4: Confusion matrix for HyperSurf.
TABLE II: Surface Recognition by HyperSurf
@@ IV-A Preparation of the stand description file
Furthermore, to accurately simulate various surfaces such as tiles, rubber, and brick within the Gazebo environment, meticulous specifications were established. These surface characteristics were detailed comprehensively in the URDF file, with each parameter meticulously tuned to replicate the physical attributes of the real-world surfaces. Key parameters, including the friction coefficient, damping coefficient, and stiffness coefficient, were precisely calibrated to ensure that the simulation dynamics closely mirror the actual conditions. The process of transferring data on surface properties from the real world to the simulation involved careful calibration of these coefficients. The goal was to align the virtual IMU data collected during simulated experiments (leg jumps) with the IMU data recorded in real-world tests. This iterative calibration process ensures that the virtual environment accurately reflects the behavior of the physical surfaces, allowing for a realistic and reliable simulation of surface interactions.
@@ IV-B Kinematics Control Implementation
Leg kinematics control was realized using symbolic and numerical computation techniques. Forward kinematics determined leg component positions, while inverse kinematics was used to adjust the joint angles to achieve specific end-effector positions. Symbolic computation was performed using SymPy, which facilitated the calculation of transformation matrices and coordinates essential for understanding the geometric relationships between different parts of the leg. CasADi was utilized for numerical optimization, handling the task of adjusting joint angles to meet the required end-effector positions with high precision. This integrated approach, combining symbolic and numerical methods, enabled fine-tuned control of the leg’s movements. It ensured effective interaction with various surfaces and environments within the Gazebo simulator, allowing for robust data collection and analysis.
@@ IV-C Dataset Collection
V Novelty of HyperSurf System
The HyperSurf system presents several innovative advancements that significantly enhance the capabilities of quadruped robot leg surface recognition, both in real-world and simulated environments. The key novelties of the system can be categorized into three primary areas: boosting the speed and efficiency of dataset collection, optimizing the workforce required for data acquisition, and enhancing the accuracy and transferability of data between physical and simulated environments.
V-A Boosting Dataset Collection Efficiency
One of the most notable innovations of the HyperSurf system is its ability to accelerate the process of dataset collection. Traditionally, collecting datasets for surface recognition involved running extensive tests with a fully operational quadruped robot, which is both time-consuming and resource-intensive. By contrast, HyperSurf introduces a one-leg test setup that significantly streamlines this process. On average, the GRU analysis speed was boosted by a factor of 7.5. This remarkable increase in efficiency is primarily due to the rapid deployment of the system and its capability to operate autonomously in controlled laboratory conditions. The reduction in complexity and the focus on a single leg not only speeds up the data collection process but also allows for more frequent iterations and refinements, leading to a faster development cycle.
V-B Optimizing Workforce for Dataset Collection
The HyperSurf system also demonstrates a significant reduction in the manpower required for dataset collection. In previous approaches, data collection was a labor-intensive process that often required the involvement of more than five team members, particularly due to the extensive duration of the experiments and the need to manage a complex, fully assembled quadruped robot. HyperSurf, however, simplifies the experimental setup by utilizing a static, one-leg configuration in a laboratory environment. This simplification not only shortens the time required for each dataset collection session but also reduces the number of personnel needed. In this study, the entire data collection process was efficiently conducted by just two researchers. This reduction in team size not only lowers operational costs but also minimizes the coordination challenges often associated with larger teams, further streamlining the research and development process.
@@ V-B Optimizing Workforce for Dataset Collection
V-C Enhancing Data Transfer and Simulation Accuracy
A critical advancement introduced by the HyperSurf system is the improved transferability of data between physical and simulated environments. By leveraging data from an IMU sensor placed on the stand, the HyperSurf system enables empirical modeling of surfaces within the Gazebo simulation environment. As shown in Fig. 5, this approach facilitates the creation of highly accurate digital twins of physical surfaces. The accuracy of this modeling is highlighted in Fig. 6, where a comparison between the angular velocities during jumps of the simulated model and the real robotic leg is presented. The results demonstrate a high degree of correlation, with the simulation achieving an accuracy of 98% relative to the real-world data.
@@ VI Conclusion and Future Work
The HyperSurf technology comprises a specialized leg test setup designed to enhance surface recognition capabilities in quadruped robots. It integrates components such as a mechanical leg with three degrees of freedom, BLDC motors for movement, and an IMU sensor for feedback collection. This setup assists the DogSurf classification model in achieving a precision of up to 98%.
Future work will focus on demonstrating how the HyperSurf surface recognition system can enhance quadruped platforms. Feedback from the GRU, processed with the DogSurf architecture, will be utilized to adapt gaits to different surfaces, adjust BLDC motor parameters, and incorporate additional mechanics to improve energy efficiency and terrain traversal. Furthermore, the development of the quadruped platform [15] will continue, integrating the system to enhance its capabilities based on surface recognition feedback. Additionally, efforts will be made to extend the system to object grasping, leveraging surface property information to improve grasping effectiveness.
In summary, the HyperSurf technology not only marks a significant milestone in surface recognition for quadruped robots but also sets the stage for broader applications across industries such as search and rescue, agriculture, and industrial inspection. This research underscores the potential of integrating advanced sensory and control systems to enable robust and adaptive autonomous systems capable of navigating complex environments with precision and reliability.
Acknowledgements
Research reported in this publication was financially supported by the RSF grant No. 24-41-02039. Thanks to Liaisan Safarova for installing the operating system on the PC and helping with the assembly!